### v1

# Surface Plasmon Polaritons: Creation Dynamics and Interference of Slow and Fast Propagating SPPs at a Temporal Boundary

Abstract
We establish the theoretical framework for a material system that supports surface plasmon polaritions (SPPs) excited by a dipole excitation, where the media configuration suddenly changes at a temporal boundary. We employ three-dimensional Green’s function analysis in the Laplace transform domain. We use this framework to demonstrate dynamic SPP formation and time-boundary-induced interference of slow and fast propagating SPPs. This analysis provides insight into how SPPs are formed in time and how they interfere at a temporal boundary.
Index Terms:
Green’s function, Laplace transform, plasmonic waveguide, surface plasmon polariton (SPP), time-varying media.
I Introduction
Electromagnetic wave propagation in time-varying media has long been a topic of research, with early investigations focusing on the electromagnetic response in media undergoing temporal changes to the material parameters. These early investigations considered the propagation of waves within a spatially homogeneous time-varying dielectric [1], pulsed excitations within a spatially homogeneous time-varying plasma [2], and a spatially homogeneous time-varying dielectric (a vacuum-dielectric half-space, where the waves are incident on the dielectric from the vacuum space) [3]. Furthermore, [4] considered waves within a spatially homogeneous dielectric, where the dielectric suddenly changes at a temporal boundary, using the Laplace transform technique, and, finally, notably, [5] considered a dipole excitation (electric dipole point source) within a spatially homogeneous dielectric (vacuum), where the dielectric suddenly changes to a plasma at a temporal boundary, using the Laplace transform technique.
In all of these cases, and subsequent research [6, 7, 8, 9, 10, 11, 12, 13], it is established that a time-varying media platform results in unique electromagnetic phenomenon not seen in time-static media, e.g., frequency shifting and frequency splitting of the incident wave at the temporal boundary, due to the conservation of momentum, which results in forward and backward waves (frequency splitting) at different frequencies than the incident wave (frequency shifting) in the temporal region after the temporal boundary.
Due to the irreversibility of time, the waves scattered at a temporal boundary can only occur in the temporal region after the temporal boundary, i.e., they can’t travel back to the past (to the temporal region before the temporal boundary). This behavior is contrary to that of wave scattering at a spatial boundary, resulting in reflected waves and transmitted waves relative to the spatial boundary.
Lately, there has been a resurgence in this research topic [7, 6, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 10, 24, 25, 26, 27, 11, 28, 29, 12, 30, 31, 32, 33, 34, 35, 36, 37, 38], driven by the potential for the applications in plasmonics (nano-photonics, nano-optics, photonic metamaterials), such as magnet-free nonreciprocity [15, 29], temporal aiming [21, 39], and extreme energy transformations [28], to name a few. There has been some work on time-varying media systems that support SPPs [16, 18, 22]; however, these deal with existing propagating SPPs encountering a temporal boundary. To the best of our knowledge, there has not been work on dipole excitation of SPPs in a time-varying system. This is the topic of this work, where such a system allows for the analysis of the interaction among dipole excitations, where we have the ability to consider the interactions at the moment of SPP creation (e.g., as a time-varying system changes from one that does not support SPPs to one that supports SPP propagation).
The use of a time-varying media platform allows for the utilization of temporal modulation, which, when combined with plasmonic waveguides (reciprocal or nonreciprocal), may allow for modifying the resonance or direction of energy propagation in the system. This then may also allow for more freedom and practicality in manipulating the electromagnetic response of plasmonic waveguides in time, which in turn may enable more efficient and tunable interactions among dipole excitations.
The typical approach to study these types of systems is to define an unbounded homogeneous space, where abrupt changes in time of the electromagnetic properties of a material in this space create temporal boundaries that replace the spatial boundaries of a related time-invariant material configuration. At these temporal boundaries momentum is conserved, however, the angular frequencies vary in time, i.e., they are specific to the temporal region before and after the switching event [6, 7, 11, 10, 12, 29], in order for the dispersion relation to be satisfied for each temporal region. That is, momentum is conserved at these temporal boundaries due to the preserved spatial homogeneity, and due to the need for the temporal boundary conditions to be satisfied everywhere in space. At the temporal boundary, all components of the electric displacement $\mathbf{D}$ and the magnetic flux density $\mathbf{B}$ are continuous, regardless of the medium model (e.g., nondispersive or dispersive), for any position in space [6, 7, 11, 10, 12, 29]. In the case of dispersive media [7], the nonlocal in time constitutive relations lead to the continuity of all of the components of $\mathbf{E}$, $\mathbf{H}$, and $\mathbf{J}$ at the temporal boundary for any position in space. The fields are then matched at the temporal boundary for all field modes allowed to propagate. That is, apply momentum conservation at the temporal boundary and then use the dispersion relation to determine the allowed modal frequencies. Then, determine the allowed field modes and match the fields at the temporal boundary to solve for the time-dependent field amplitudes. We then have the field solutions for each temporal region at the temporal boundary. This is the approach used in most cases where the electromagnetic waves can be decomposed (modal expansions can be performed) to obtain the wavenumbers for the propagating waves in each temporal region.
However, this method becomes difficult in the case of a time-varying media system that supports SPPs excited by a dipole excitation (electric dipole point source), e.g., a single dielectric-plasma interface with a dipole excitation above the interface, because there are numerous wavenumbers, in multiple directions, supporting multiple propagating waves (radiation modes (radiating waves), SPP modes (surface waves), Brewster (bulk) modes (waves in the plasma (bulk) region)) that satisfy the propagation constraints of the spatial configuration. Additionally, for each one of these allowed propagation constants there are numerous modal temporal frequencies that satisfy the propagation constraints of the temporal configuration (momentum conservation at the temporal boundary).
Therefore, we instead utilize the Laplace transform technique, which accounts for the temporal boundary conditions in the initial conditions of the Laplace transform. The inverse Laplace transform naturally sums up all of the contributions from the allowed propagating modes for the allowed modal temporal frequencies ($s$-values) for the allowed modal spatial frequencies ($q$-values).
In the subsequent sections we establish the theoretical framework for the specific configuration of a time-varying media system that supports SPPs, in at least one temporal region, or both, excited by a dipole source, where the media configuration suddenly changes at a temporal boundary. We first state Maxwell’s equations in the Laplace transform domain, derive the inhomogenous electric field wave equation in the Laplace transform domain for generic media, i.e., linear, inhomogeneous, anisotropic, temporally dispersive (temporally nonlocal), and spatially nondispersive (local) media, and then illustrate the use of Green’s functions to solve the inhomogenous wave equation to obtain the electric field.
We then consider the special case of isotropic media and one temporal boundary, where we derive the corresponding Green’s functions and demonstrate dynamic SPP formation and interference of slow and fast propagating SPPs at the temporal boundary.
The model for this configuration, seen in Fig. 1, defines two temporal regions (one temporal boundary).
Figure 1: Model for the general case of a single interface between two different materials ($\mu_{2}$,$\varepsilon_{2}$ for $z>0$, $\mu_{1}$,$\varepsilon_{1}$ for $z<0$) with an electric dipole point source, $\widehat{\mathbf{z}}j_{zs}$, located along the interface surface, where we consider two temporal regions (one temporal boundary). The double subscript notation is described in the text. Additionally, $q_{x}$, $q_{y}$, and $\phi$ define the coordinate system in the momentum space.
We use dispersive materials in the model, in order to ensure that the temporal boundary conditions are met for $\mathbf{E}$, $\mathbf{H}$, and $\mathbf{J}$ (i.e., we have a temporally causal system).
Note that in this work we simply consider that the media parameters are suddenly changed. In practice, this can be achieved in material systems with free electrons, such as plasmas, semiconductors, and two-dimensional systems such as graphene, by the change in a DC bias (DC magnetic field or DC voltage). Although those systems are anisotropic, here we consider isotropic materials for simplicity; the extension to anisotropic systems is straightforward as the system is described by the Green’s function.
II Theoretical Model
II-A Maxwell’s Equations in the Laplace Transform Domain
Maxwell’s equations in matter, in the Laplace transform domain, can be written as [40, 41]
×$\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}(\mathbf{r},s)$ $=-s\mu_{0}\underline{\boldsymbol{\mu}}_{r}(\mathbf{r},s)\cdot\mathbf{H}(\mathbf{r},s)$ (1)
$+\mu_{0}\mathbf{H}(\mathbf{r},t^{-})+\mu_{0}\mathbf{M}(\mathbf{r},t^{-}),$
×$\mathbf{\boldsymbol{\nabla}}\times\mathbf{H}(\mathbf{r},s)$ $=s\varepsilon_{0}\underline{\boldsymbol{\varepsilon}}_{r}(\mathbf{r},s)\cdot\mathbf{E}(\mathbf{r},s)$ (2)
$-\varepsilon_{0}\mathbf{E}(\mathbf{r},t^{-})-\mathbf{P}(\mathbf{r},t^{-})+\mathbf{j}_{s}(\mathbf{r},s),$
$\mathbf{\boldsymbol{\nabla}}\cdot\mathbf{D}(\mathbf{r},s)$ $=\rho_{s}(\mathbf{r},s),$ (3)
$\mathbf{\boldsymbol{\nabla}}\cdot\mathbf{B}(\mathbf{r},s)$ $=0,$ (4)
where $\mathbf{E}$ is the electric field intensity, $\mathbf{B}$ is the magnetic flux density, $\rho$ is the source charge density, $\mathbf{j}$ is the source current density, $\varepsilon_{0}$ is the permittivity of vacuum, $\mu_{0}$ is the permeability of vacuum, $\underline{\boldsymbol{\varepsilon}}_{r}(\mathbf{r},s)$ is the relative permittivity tensor, $\underline{\boldsymbol{\mu}}_{r}(\mathbf{r},s)$ is the relative permeability tensor, $\mathbf{D}$ is the electric flux density, $\mathbf{H}$ is the magnetic field intensity, $\mathbf{P}$ is the polarization, and $\mathbf{M}$ is the magnetization. Additionally, variables that are a function of $t^{-}$ are the initial conditions for that variable at the time just before time $t$.
II-B Inhomogeneous Wave Equation for the Electric Field in the Laplace Transform Domain
We now derive the inhomogeneous wave equation for the electric field in the Laplace transform domain. We can then solve for the electric field in each medium. Assuming non-magnetic media, $\underline{\boldsymbol{\mu}}_{r}(\mathbf{r},s)=\underline{\mathbf{I}}$ and $\mathbf{M}(\mathbf{r},t^{-})=\mathbf{0}$, multiplying both sides of (1) by $\mu_{\mathrm{0}}^{-1}\underline{\boldsymbol{\mu}}^{-1}(\mathbf{r},s)$, taking the curl of both sides, inserting (2), and multiplying both sides by $\mu_{0}$, we have
××$\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}(\mathbf{r},s)$ $=-\frac{s^{2}}{c^{2}}\underline{\boldsymbol{\varepsilon}}_{r}(\mathbf{r},s)\cdot\mathbf{E}(\mathbf{r},s)$ (5)
$-s\mu_{0}\mathbf{j}_{s}(\mathbf{r},s)+\frac{s}{c^{2}}\mathbf{E}(\mathbf{r},t^{-})$
$+\mu_{0}s\mathbf{P}(\mathbf{r},t^{-})+\mu_{0}\mathbf{\boldsymbol{\nabla}}\times\mathbf{H}(\mathbf{r},t^{-}),$
where we used $\nicefrac{{1}}{{c^{2}}}=\mu_{0}\varepsilon_{0}$. Equation (5) is the inhomogenous wave equation for the electric field in the Laplace transform domain.
In order to obtain a more usable form, where we can incorporate common initial conditions, using Ampere’s law in the time domain, ×$\mathbf{\boldsymbol{\nabla}}\times\mathbf{H}(\mathbf{r},t)=\frac{\partial}{\partial t}\mathbf{D}(\mathbf{r},t)+\mathbf{j}_{s}(\mathbf{r},t)\rightarrow\mathbf{\boldsymbol{\nabla}}\times\mathbf{H}(\mathbf{r},t^{-})=\frac{\partial}{\partial t}\mathbf{D}(\mathbf{r},t)\left|{}_{t=t^{-}}\right.+\mathbf{j}_{s}(\mathbf{r},t^{-})$, and the electric constitutive relation in the time domain, $\mathbf{D}(\mathbf{r},t)=\varepsilon_{0}\mathbf{E}(\mathbf{r},t)+\mathbf{P}(\mathbf{r},t)\rightarrow\mathbf{P}(\mathbf{r},t^{-})=\mathbf{D}(\mathbf{r},t^{-})-\varepsilon_{0}\mathbf{E}(\mathbf{r},t^{-})$, we can write (5) as
××$\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}(\mathbf{r},s)$ $=-\frac{s^{2}}{c^{2}}\underline{\boldsymbol{\varepsilon}}_{r}(\mathbf{r},s)\cdot\mathbf{E}(\mathbf{r},s)$ (6)
$-s\mu_{0}\mathbf{j}_{s}(\mathbf{r},s)+\mu_{0}\mathbf{j}_{s}(\mathbf{r},t)\left|{}_{t=t^{-}}\right.$
$+\mu_{0}\frac{\partial}{\partial t}\mathbf{D}(\mathbf{r},t)\left|{}_{t=t^{-}}\right.+\mu_{0}s\mathbf{D}(\mathbf{r},t)\left|{}_{t=t^{-}}\right..$
We can use (6) to solve for the electric field in any temporal region, where the initial conditions are known at the temporal boundary immediately preceding the temporal region. We assume that there is no change in the induced current density just after the time-change, i.e., at $t=t_{\alpha}^{+}$, where $\alpha=0,1,2,...$. That is, although the material permittivity is modeled as changing value instantaneously, the material response (electrons) cannot change instantaneously due to their mass. This leads to the temporal boundary conditions (BCs); all components of the electric field $\mathbf{E}$, the magnetic field $\mathbf{H}$, and the current density $\mathbf{J}$ are continuous at the temporal boundary. Then the fields are the same just before and just after the time-change with respect to the media (and consequently at the time-change $t=t_{\alpha}$), i.e., at $t=t_{\alpha}^{-}$ and $t=t_{\alpha}^{+}$. In other words, we can use $t=t_{\alpha}^{-}$, $t=t_{\alpha}$, or $t=t_{\alpha}^{+}$ for the initial condition parameters.
We can use (6) as is for a known initial condition in the time domain, which we will set for the case of temporal region 1 since we set the initial conditions that start the system. Once we solve for the fields in temporal region 1 we will have the initial conditions to solve for the fields in temporal region 2.
II-C One Temporal Boundary
Here, we consider the case of one temporal boundary at $t=t_{0}=0$. We assume the sources and fields in temporal region 1 are time-harmonic (monochromatic, sinusoidal steady-state) with time variations of the form $e^{-i\omega_{0}t}$. At this point we formally adopt the notation of $x_{mn}$, where $x$ is some parameter, where $m=1,2,3,...$ designates its spatial region and $n=1,2,3,...$ designates its temporal region. We then write (6) for temporal region 2 as
××$\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}_{m2}(\mathbf{r},s)$ $=-\frac{s^{2}}{c^{2}}\underline{\boldsymbol{\varepsilon}}_{rm2}(\mathbf{r},s)\cdot\mathbf{E}_{m2}(\mathbf{r},s)$ (7)
$-s\mu_{0}\mathbf{j}_{s22}(\mathbf{r},s)+\mu_{0}\mathbf{j}_{s21}(\mathbf{r},t)\left|{}_{t_{0}=0}\right.$
$+\mu_{0}\frac{\partial}{\partial t}\mathbf{D}_{m1}(\mathbf{r},t)\left|{}_{t_{0}=0}\right.$
$+\mu_{0}s\mathbf{D}_{m1}(\mathbf{r},t)\left|{}_{t_{0}=0}\right.,$
where $\mathbf{j}_{s21}(\mathbf{r},t)=\mathrm{Re}\left\{\mathbf{j}_{s21}(\mathbf{r})e^{-i\omega_{0}t}\right\}\rightarrow\mathbf{j}_{s21}(\mathbf{r})e^{-i\omega_{0}t}$ and $\mathbf{D}_{m1}(\mathbf{r},t)=\mathrm{Re}\left\{\mathbf{D}_{m1}(\mathbf{r})e^{-i\omega_{0}t}\right\}\rightarrow\mathbf{D}_{m1}(\mathbf{r})e^{-i\omega_{0}t}$. [^1] We drop the $\mathrm{Re}\left\{\cdot\right\}$ notation to simplify calculations, where we will then take the $\mathrm{Re}\left\{\cdot\right\}$ of the final expressions in the time domain to ensure that the values are real in the time domain. We note that this only applies to the expressions in temporal region 1, where we assumed time-harmonic conditions. We assume that the source current density remains the same for all of time in all temporal regions, then we can write $\mathbf{j}_{s21}(\mathbf{r},t)\rightarrow\mathbf{j}_{s}(\mathbf{r})e^{-i\omega_{0}t}$ and $\mathbf{j}_{s22}(\mathbf{r},s)=\mathbf{j}_{s21}(\mathbf{r},s)=\mathcal{L}\left\{\mathbf{j}_{s21}(\mathbf{r},t)\right\}=\mathcal{L}\left\{\mathbf{j}_{s}(\mathbf{r})e^{-i\omega_{0}t}\right\}=\frac{1}{s+i\omega_{0}}\mathbf{j}_{s}(\mathbf{r})$. Additionally, with $\mathbf{D}(\mathbf{r})=\varepsilon_{0}\underline{\boldsymbol{\varepsilon}}_{r}(\mathbf{r},\omega_{0})\cdot\mathbf{E}(\mathbf{r})$, we can write $\mathbf{D}_{m1}(\mathbf{r})=\varepsilon_{0}\underline{\boldsymbol{\varepsilon}}_{rm1}(\mathbf{r},\omega_{0})\cdot\mathbf{E}_{m1}(\mathbf{r})$. We can then can write (7) as
××$\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}_{m2}(\mathbf{r},s)$ $=-\frac{s^{2}}{c^{2}}\underline{\boldsymbol{\varepsilon}}_{rm2}(\mathbf{r},s)\cdot\mathbf{E}_{m2}(\mathbf{r},s)$ (8)
$+\frac{i\omega_{0}\mu_{0}}{s+i\omega_{0}}\mathbf{j}_{s}(\mathbf{r})$
$+\frac{s-i\omega_{0}}{c^{2}}\underline{\boldsymbol{\varepsilon}}_{rm1}(\mathbf{r},\omega_{0})\cdot\mathbf{E}_{m1}(\mathbf{r}),$
which is the inhomogenous wave equation for the electric field in the Laplace transform domain (when used for spatial region 1 we set $\mathbf{j}_{s}(\mathbf{r})$ to zero).
We can use the inhomogenous wave equation for the electric field in the spatial regions in temporal region 1,
××$\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}_{m1}(\mathbf{r})$ $=\frac{\omega_{0}^{2}}{c^{2}}\underline{\boldsymbol{\varepsilon}}_{rm1}(\mathbf{r},\omega_{0})\cdot\mathbf{E}_{m1}(\mathbf{r})$ (9)
$+i\omega_{0}\mu_{0}\mathbf{j}_{s}(\mathbf{r}),$
to solve for the electric field in the spatial regions in temporal region 1 (when used for spatial region 1 we set $\mathbf{j}_{s}(\mathbf{r})$ to zero). We can solve for the magnetic field in temporal region 1 using
$\mathbf{H}_{m1}(\mathbf{r})$ $=\frac{1}{i\omega_{0}\mu_{0}}\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}_{m1}(\mathbf{r}).$ (10)
II-D Obtaining the Standard Form for the Wave Equation
The solution for the fields in temporal region 1 follows any usual (non-time-varying media) method, such as using Green’s functions, which then leads to the field values for the temporal BCs. For temporal region 2 we also use the Green’s function, although we first need to cast (8) into a standard form (from which the Green’s function follows in a straightforward manner). In (8) we can see that we have an extra term on the right hand side of the wave equation (a source-like term arising from the initial condition), which results in the wave equation not being in the standard form. Following [5], we can obtain the standard form for the wave equation by performing a transformation regarding this term. We start by rewriting $\mathbf{E}_{m1}(\mathbf{r})$ as
$\mathbf{E}_{m1}(\mathbf{r})$ $=\mathbf{E}_{m1}(\mathbf{r})\cdot\left(\underline{\mathbf{F}}_{1}(\mathbf{r},s)+\underline{\mathbf{F}}_{2}(\mathbf{r},s)\right),$ (11)
where we assume $\underline{\mathbf{F}}_{1}$ and $\underline{\mathbf{F}}_{2}$ are tensors, functions of $\mathbf{r}$ and $s$, and they satisfy the unity relation $\underline{\mathbf{F}}_{1}(\mathbf{r},s)+\underline{\mathbf{F}}_{2}(\mathbf{r},s)=\underline{\mathbf{I}}$. Using (9), we can write
$\mathbf{E}_{m1}(\mathbf{r})$ $=\frac{\underline{\boldsymbol{\varepsilon}}_{rm1}^{-1}(\mathbf{r},\omega_{0})}{k_{0}^{2}}\cdot\left(\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}_{m1}(\mathbf{r})-i\omega_{0}\mu_{0}\mathbf{j}_{s}(\mathbf{r})\right),$ (12)
where we note that $k_{0}=\nicefrac{{\omega_{0}}}{{c}}$. Then (11) can be written as
$\mathbf{E}_{m1}(\mathbf{r})$ $=\mathbf{E}_{m1}(\mathbf{r})\cdot\underline{\mathbf{F}}_{1}(\mathbf{r},s)$ (13)
$+\frac{\underline{\boldsymbol{\varepsilon}}_{rm1}^{-1}(\mathbf{r},\omega_{0})}{k_{0}^{2}}\cdot\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}_{m1}(\mathbf{r})\cdot\underline{\mathbf{F}}_{2}(\mathbf{r},s)$
$-\frac{\underline{\boldsymbol{\varepsilon}}_{rm1}^{-1}(\mathbf{r},\omega_{0})}{k_{0}^{2}}\cdot i\omega_{0}\mu_{0}\mathbf{j}_{s}(\mathbf{r})\cdot\underline{\mathbf{F}}_{2}(\mathbf{r},s).$
Then, after substituting (13) into (8) and using the resulting equation and the unity relation to solve for $\underline{\mathbf{F}}_{1}$ and $\underline{\mathbf{F}}_{2}$, we can write
××$\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}_{m2}^{\prime}(\mathbf{r},s)$ $=-\frac{s^{2}}{c^{2}}\underline{\boldsymbol{\varepsilon}}_{rm2}(\mathbf{r},s)\cdot\mathbf{E}_{m2}^{\prime}(\mathbf{r},s)-s\mu_{0}\mathbf{j}_{s}^{\prime}(\mathbf{r},s),$ (14)
where
$\mathbf{E}_{m2}^{\prime}(\mathbf{r},s)$ $\equiv\mathbf{E}_{m2}(\mathbf{r},s)-\left(s-i\omega_{0}\right)\left(s^{2}\underline{\boldsymbol{\varepsilon}}_{rm1}^{-1}(\mathbf{r},\omega_{0})\cdot\underline{\boldsymbol{\varepsilon}}_{rm2}(\mathbf{r},s)+\omega_{0}^{2}\underline{\mathbf{I}}\right)^{-1}\cdot\mathbf{E}_{m1}(\mathbf{r}),$ (15)
$\mathbf{j}_{s}^{\prime}(\mathbf{r},s)$ $\equiv\left(\frac{-i\omega_{0}}{s\left(s+i\omega_{0}\right)}-\frac{-i\omega_{0}}{s}\left(s-i\omega_{0}\right)\left(s^{2}\underline{\boldsymbol{\varepsilon}}_{rm1}^{-1}(\mathbf{r},\omega_{0})\cdot\underline{\boldsymbol{\varepsilon}}_{rm2}(\mathbf{r},s)+\omega_{0}^{2}\underline{\mathbf{I}}\right)^{-1}\right)\cdot\mathbf{j}_{s}(\mathbf{r}).$ (16)
Equation (14) is now in the standard form for a wave equation, which aligns with (9).
We can determine the magnetic field, with $\mathbf{H}_{m1}(\mathbf{r},t^{-})\rightarrow\mathbf{H}_{m1}(\mathbf{r},t)\left|{}_{t_{0}=0}\right.$, where $\mathbf{H}_{m1}(\mathbf{r},t)=\mathrm{Re}\left\{\mathbf{H}_{m1}(\mathbf{r})e^{-i\omega_{0}t}\right\}\rightarrow\mathbf{H}_{m1}(\mathbf{r})e^{-i\omega_{0}t}$, as
$\mathbf{H}_{m2}(\mathbf{r},s)$ $=\frac{1}{-s\mu_{0}}\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}_{m2}(\mathbf{r},s)+\frac{1}{i\omega_{0}\mu_{0}s}\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}_{m1}(\mathbf{r}).$ (17)
Using (15) we can write an expression for $\mathbf{E}_{m2}(\mathbf{r},s)$ and substitute it into (17) to obtain
$\mathbf{H}_{m2}(\mathbf{r},s)$ $=\mathbf{H}_{m2}^{\prime}(\mathbf{r},s)+\frac{1}{-s\mu_{0}}\mathbf{\boldsymbol{\nabla}}\times\left(\left(s-i\omega_{0}\right)\left(s^{2}\underline{\boldsymbol{\varepsilon}}_{rm1}^{-1}(\mathbf{r},\omega_{0})\cdot\underline{\boldsymbol{\varepsilon}}_{rm2}(\mathbf{r},s)+\omega_{0}^{2}\underline{\mathbf{I}}\right)^{-1}\cdot\mathbf{E}_{m1}(\mathbf{r})\right)+\frac{1}{i\omega_{0}\mu_{0}s}\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}_{m1}(\mathbf{r}),$ (18)
where $\mathbf{H}_{m2}^{\prime}(\mathbf{r},s)=\frac{1}{-s\mu_{0}}\mathbf{\boldsymbol{\nabla}}\times\mathbf{E}_{m2}^{\prime}(\mathbf{r},s)$.
In the case of piece-wise homogeneous layers, $\underline{\boldsymbol{\varepsilon}}_{rmn}(\mathbf{r})\rightarrow\underline{\boldsymbol{\varepsilon}}_{rmn}$, and isotropic media, $\underline{\boldsymbol{\varepsilon}}_{rmn}\rightarrow\underline{\mathbf{I}}\varepsilon_{rmn}$, we can write $\mathbf{E}_{m2}(\mathbf{r},s)=\mathbf{E}_{m2}^{\prime}(\mathbf{r},s)+A_{m2}^{e}(s)\mathbf{E}_{m1}(\mathbf{r})$, $\mathbf{H}_{m2}(\mathbf{r},s)=\mathbf{H}_{m2}^{\prime}(\mathbf{r},s)+A_{m2}^{m}(s)\mathbf{H}_{m1}(\mathbf{r})$, and $\mathbf{j}_{s}^{\prime}(\mathbf{r},s)=A_{22}^{s}(s)\mathbf{j}_{s}(\mathbf{r})$, where
$A_{m2}^{e}(s)$ $=\frac{\left(s-i\omega_{0}\right)\varepsilon_{rm1}(\omega_{0})}{s^{2}\varepsilon_{rm2}(s)+\omega_{0}^{2}\varepsilon_{rm1}(\omega_{0})},$ (19)
$A_{m2}^{m}(s)$ $=\frac{s\varepsilon_{rm2}(s)-i\omega_{0}\varepsilon_{rm1}(\omega_{0})}{s^{2}\varepsilon_{rm2}(s)+\omega_{0}^{2}\varepsilon_{rm1}(\omega_{0})},$ (20)
and
$A_{22}^{s}(s)=A_{2}^{s}(s)$ $=\frac{-i\omega_{0}s^{2}\left(\varepsilon_{r22}(s)-\varepsilon_{r21}(\omega_{0})\right)}{s\left(s+i\omega_{0}\right)\left(s^{2}\varepsilon_{r22}(s)+\omega_{0}^{2}\varepsilon_{r21}(\omega_{0})\right)}.$ (21)
II-E Green’s Functions in the Laplace Transform Domain
We can solve (14) using the Green’s tensor that satisfies
(××$\left(\mathbf{\boldsymbol{\nabla}}\times\mathbf{\boldsymbol{\nabla}}\times+\frac{s^{2}}{c^{2}}\underline{\boldsymbol{\varepsilon}}_{rm2}(\mathbf{r},s)\cdot\right)\underline{\mathbf{G}}_{m2}^{\prime}(\mathbf{r},\mathbf{r}^{\prime},s)$ $=\underline{\mathbf{I}}\delta(\mathbf{r}-\mathbf{r}^{\prime}),$ (22)
where $\underline{\mathbf{G}}_{m2}^{\prime}(\mathbf{r},\mathbf{r}^{\prime},s)$ is the electric field Green’s function tensor for the primed electric field in the Laplace transform domain.
II-E1 Spatial Boundary Conditions for the Fields
We need to enforce the spatial boundary conditions for the primed-fields previously defined at the interface. The well established spatial boundary conditions for the fields at an interface between two media, where there are no charges or sources along the interface (i.e., the two media are not perfect conductors and there are no impressed sources placed along the interface), are [42]
$E_{x2n}-E_{x1n}=0,$ $\;E_{y2n}-E_{y1n}=0,$ (23)
$H_{x2n}-H_{x1n}=0,$ $\;H_{y2n}-H_{y1n}=0.$ (24)
In the case of temporal region 2, from (15) and (18) the spatial boundary conditions on $\mathbf{E}_{m2}(\mathbf{r},s)$ and $\mathbf{H}_{m2}(\mathbf{r},s)$ translate into spatial boundary conditions on $\mathbf{E}_{m2}^{\prime}(\mathbf{r},s)$ and $\mathbf{H}_{m2}^{\prime}(\mathbf{r},s)$, where, for piece-wise homogeneous layers and isotropic media, and $\beta=x,y$, they are
$E_{\beta 22}^{\prime}-E_{\beta 12}^{\prime}$ $=A_{12}^{e}(s)E_{\beta 11}-A_{22}^{e}(s)E_{\beta 21},$ (25)
$H_{\beta 22}^{\prime}-H_{\beta 12}^{\prime}$ $=A_{12}^{m}(s)H_{\beta 11}-A_{22}^{m}(s)H_{\beta 21}.$ (26)
II-E2 Isotropic Media
We now establish the Green’s function solutions that satisfy (22), consistent with (25) and (26), for the case of piece-wise homogeneous layers and isotropic media. This framework can be extended to obtain solutions for the case of anisotropic media by following well-known methods for obtaining Green’s functions for more complicated materials. We derive the Green’s functions and the corresponding electric field solutions for the complete structure (Fig. 1) utilizing the Hertz potentials [43, 44], where we define a vertical point dipole current source as a unit point current source $\mathbf{j}_{s}(\mathbf{r},t)=\widehat{\mathbf{z}}\delta(\mathbf{r}-\mathbf{r}_{0})\cos\left(-\omega_{0}t\right)=\mathrm{Re}\left\{\mathbf{j}_{s}(\mathbf{r})e^{-i\omega_{0}t}\right\}$, where $\mathbf{j}_{s}(\mathbf{r})=\widehat{\mathbf{z}}\delta(\mathbf{r}-\mathbf{r}_{0})$. We note that the Green’s functions here for temporal region 2 are new, since the $\mathbf{E}^{\prime}$ and $\mathbf{H}^{\prime}$ fields satisfy different spatial boundary conditions than the usual $\mathbf{E}$ and $\mathbf{H}$ fields.
The Hertz potential Green’s function components are
[…]
$\frac{e^{-p_{2n}\left|z-z^{\prime}\right|}}{2p_{2n}}e^{i\left(q_{x}\left(x-x^{\prime}\right)+q_{y}\left(y-y^{\prime}\right)\right)}$
$=\frac{e^{ik_{2n}\sqrt{\rho^{2}+\left(z-z^{\prime}\right)^{2}}}}{4\pi\sqrt{\rho^{2}+\left(z-z^{\prime}\right)^{2}}},$
[…]
$R_{2n}e^{-p_{2n}z}\frac{e^{-p_{2n}z^{\prime}}}{2p_{2n}}e^{i\left(q_{x}\left(x-x^{\prime}\right)+q_{y}\left(y-y^{\prime}\right)\right)},$
[…]
$T_{1n}e^{p_{1n}z}\frac{e^{-p_{2n}z^{\prime}}}{2p_{2n}}e^{i\left(q_{x}\left(x-x^{\prime}\right)+q_{y}\left(y-y^{\prime}\right)\right)},$
where we can write $g_{zz2n}(\mathbf{r},\mathbf{r}^{\prime})=g_{zz2n}^{p}(\mathbf{r},\mathbf{r}^{\prime})+g_{zz2n}^{s}(\mathbf{r},\mathbf{r}^{\prime})$, with the $p$ superscript designating the principal field and the $s$ superscript designating the scattered field, $\rho=\sqrt{\left(x-x^{\prime}\right)^{2}+\left(y-y^{\prime}\right)^{2}}$, and where
$R_{21}(\omega_{0})=\frac{\frac{\varepsilon_{11}}{\varepsilon_{21}}p_{21}-p_{11}}{\frac{\varepsilon_{11}}{\varepsilon_{21}}p_{21}+p_{11}},$ $\;T_{11}(\omega_{0})=\frac{2p_{21}}{\frac{\varepsilon_{11}}{\varepsilon_{21}}p_{21}+p_{11}},$ (30)
and
$R_{22}(z^{\prime},s)$ $=\frac{\frac{\varepsilon_{12}(s)}{\varepsilon_{22}(s)}p_{22}-p_{12}}{\frac{\varepsilon_{12}(s)}{\varepsilon_{22}(s)}p_{22}+p_{12}}$ (31)
$+\frac{1}{\frac{e^{-p_{22}z^{\prime}}}{2p_{22}}}$ $\frac{2p_{21}\varepsilon_{r22}(s)\left(\varepsilon_{r12}(s)\varepsilon_{r21}(\omega_{0})-\varepsilon_{r11}(\omega_{0})\varepsilon_{r22}(s)\right)}{\omega_{0}^{2}\left(p_{21}\varepsilon_{r11}(\omega_{0})+p_{11}\varepsilon_{r21}(\omega_{0})\right)\left(\varepsilon_{r21}(\omega_{0})-\varepsilon_{r22}(s)\right)}\frac{\left(s^{2}+\omega_{0}^{2}\right)\left(s^{2}p_{11}\varepsilon_{r12}(s)+\omega_{0}^{2}p_{12}\varepsilon_{r11}(\omega_{0})\right)}{\left(p_{22}\varepsilon_{r12}(s)+p_{12}\varepsilon_{r22}(s)\right)\left(s^{2}\varepsilon_{r12}(s)+\omega_{0}^{2}\varepsilon_{r11}(\omega_{0})\right)}\frac{e^{-p_{21}z^{\prime}}}{2p_{21}},$
$T_{12}(z^{\prime},s)$ $=\frac{2p_{22}}{\frac{\varepsilon_{12}(s)}{\varepsilon_{22}(s)}p_{22}+p_{12}}$ (32)
$+\frac{1}{\frac{e^{-p_{22}z^{\prime}}}{2p_{22}}}$ $\frac{2p_{21}\varepsilon_{r22}(s)\left(\varepsilon_{r12}(s)\varepsilon_{r21}(\omega_{0})-\varepsilon_{r11}(\omega_{0})\varepsilon_{r22}(s)\right)}{\omega_{0}^{2}\left(p_{21}\varepsilon_{r11}(\omega_{0})+p_{11}\varepsilon_{r21}(\omega_{0})\right)\left(\varepsilon_{r21}(\omega_{0})-\varepsilon_{r22}(s)\right)}\frac{\left(s^{2}+\omega_{0}^{2}\right)\left(s^{2}p_{11}\varepsilon_{r22}(s)-\omega_{0}^{2}p_{22}\varepsilon_{r11}(\omega_{0})\right)}{\left(p_{22}\varepsilon_{r12}(s)+p_{12}\varepsilon_{r22}(s)\right)\left(s^{2}\varepsilon_{r12}(s)+\omega_{0}^{2}\varepsilon_{r11}(\omega_{0})\right)}\frac{e^{-p_{21}z^{\prime}}}{2p_{21}},$
which are the scattering (reflection) coefficient and the transmission coefficient, respectively, where $q=\sqrt{q_{x}^{2}+q_{y}^{2}}$, $p_{m2}=\sqrt{q^{2}-k_{m2}^{2}(s)}$, $k_{m2}(s)=is\sqrt{\mu_{0}\varepsilon_{m2}(s)}$, $\varepsilon_{m2}(s)=\varepsilon_{0}\varepsilon_{rm2}(s)$ and $p_{m1}=\sqrt{q^{2}-k_{m1}^{2}(\omega_{0})}$, $k_{m1}(\omega_{0})=\omega_{0}\sqrt{\mu_{0}\varepsilon_{m1}(\omega_{0})}$, $\varepsilon_{m1}(\omega_{0})=\varepsilon_{0}\varepsilon_{rm1}(\omega_{0})$.
Note that in the limiting case of no material change at the temporal boundary $t=0$, $\varepsilon_{rm1}(\omega_{0})=\varepsilon_{rm2}(\omega_{0})\Leftrightarrow\varepsilon_{rm2}(s)$, which leads to $p_{m1}(\omega_{0})=p_{m2}(\omega_{0})\Leftrightarrow p_{m2}(s)$. Then $R_{22}(z^{\prime},s)\rightarrow R_{22}(\omega_{0})=R_{21}(\omega_{0})$ and $T_{12}(z^{\prime},s)\rightarrow T_{12}(\omega_{0})=T_{11}(\omega_{0}),$ which is the usual case for no temporal boundaries.
II-E3 Dispersion Relation
The dispersion relations for the TM surface waves (the SPPs) supported by the interface are obtained by setting the denominators of the coefficients in (31) equal to zero. That is,
$\frac{\varepsilon_{12}(s)}{\varepsilon_{22}(s)}p_{22}+p_{12}$ $=0,$ (33)
and
$\omega_{0}^{2}\left(p_{21}\varepsilon_{r11}(\omega_{0})+p_{11}\varepsilon_{r21}(\omega_{0})\right)\left(\varepsilon_{r21}(\omega_{0})-\varepsilon_{r22}(s)\right)\left(p_{22}\varepsilon_{r12}(s)+p_{12}\varepsilon_{r22}(s)\right)\left(s^{2}\varepsilon_{r12}(s)+\omega_{0}^{2}\varepsilon_{r11}(\omega_{0})\right)$ $=0.$ (34)
Solving these for $q$ we obtain
$q(s)$ $=k_{0}\sqrt{\frac{\varepsilon_{r12}(s)\varepsilon_{r22}(s)}{\varepsilon_{r12}(s)+\varepsilon_{r22}(s)}},$ (35)
for (33), and (35) and
$q(\omega_{0})$ $=k_{0}\sqrt{\frac{\varepsilon_{r11}(\omega_{0})\varepsilon_{r21}(\omega_{0})}{\varepsilon_{r11}(\omega_{0})+\varepsilon_{r21}(\omega_{0})}}$ (36)
for (34), where we see that we have SPP modes from temporal region 1 propagating in temporal region 2, in addition to those created at the temporal boundary $t=0$ and propagating in temporal region 2. Again, in the limiting case of no material change at the temporal boundary $t=0$, $R_{22}(z^{\prime},s)\rightarrow R_{22}(\omega_{0})=R_{21}(\omega_{0})$, resulting in (36), which is the usual case for no temporal boundaries, where we now have just the usual SPP modes propagating.
II-E4 Fields
[…]
$\mathbf{E}_{m1}(\mathbf{r}),$ $\>\mathbf{E}_{m2}^{\prime}(\mathbf{r},s)=C_{mn}^{e}\left(\begin{array}[]{c}\widehat{\mathbf{z}}g_{zzmn}(\mathbf{r},\mathbf{r}^{\prime})+\frac{1}{k_{mn}^{2}}\left(\widehat{\mathbf{x}}\frac{\partial}{\partial x}\frac{\partial}{\partial z}g_{zzmn}(\mathbf{r},\mathbf{r}^{\prime})+\widehat{\mathbf{y}}\frac{\partial}{\partial y}\frac{\partial}{\partial z}g_{zzmn}(\mathbf{r},\mathbf{r}^{\prime})+\widehat{\mathbf{z}}\frac{\partial^{2}}{\partial z^{2}}g_{zzmn}(\mathbf{r},\mathbf{r}^{\prime})\right)\end{array}\right),$ (38)
$\mathbf{H}_{m1}(\mathbf{r}),$ $\>\mathbf{H}_{m2}^{\prime}(\mathbf{r},s)=C_{mn}^{m}\left(\widehat{\mathbf{x}}\frac{\partial}{\partial y}g_{zzmn}(\mathbf{r},\mathbf{r}^{\prime})-\widehat{\mathbf{y}}\frac{\partial}{\partial x}g_{zzmn}(\mathbf{r},\mathbf{r}^{\prime})\right),$ (39)
where $C_{11}^{e}=i\omega_{0}\mu_{0}\left(\nicefrac{{\varepsilon_{11}}}{{\varepsilon_{21}}}\right)$, $C_{21}^{e}=i\omega_{0}\mu_{0}$, $C_{12}^{e}=A_{2}^{s}(s)\left(-s\mu_{0}\right)\left(\nicefrac{{\varepsilon_{12}(s)}}{{\varepsilon_{22}(s)}}\right)$, $C_{22}^{e}=A_{2}^{s}(s)\left(-s\mu_{0}\right)$, $C_{11}^{m}=\nicefrac{{\varepsilon_{11}}}{{\varepsilon_{21}}}$, $C_{21}^{m}=1$, $C_{12}^{m}=A_{2}^{s}(s)\left(\nicefrac{{\varepsilon_{12}(s)}}{{\varepsilon_{22}(s)}}\right)$, and $C_{22}^{m}=A_{2}^{s}(s)$
II-F Fields in the Time Domain
We can obtain the fields in the time domain in temporal region 1 using simply $\mathrm{Re}\left\{\left(\cdot\right)e^{-i\omega_{0}t}\right\}$ and in temporal region 2 using inverse Laplace transforms.
III Results
In the following we obtain results for material configurations that give a strong SPP response. In all cases the results are obtained for the configuration in Fig. 1, where, as before, we assume nonmagnetic materials, piece-wise homogeneous layers, and isotropic media.
The permittivity expression used for the dielectric regions (for fields with time variations of the form $e^{-i\omega t}$) is
$\varepsilon_{r}(\omega)$ $=\varepsilon_{r}^{\prime}+i\frac{\sigma}{\omega\varepsilon_{0}},$ (40)
which is the complex permittivity.
The permittivity expression used for the plasma regions is
$\varepsilon_{r}(\omega)=1-\frac{\omega_{p}^{2}}{\omega^{2}+i\omega\Gamma_{d}},$ (41)
which is the Drude dispersion model for the plasma, where $\omega_{p}$ is the plasma frequency (which is a function of the material free electron density) and $\Gamma_{d}$ is the damping constant (damping frequency, electron mean collision frequency, i.e., $\Gamma_{d}=\nicefrac{{1}}{{\tau}}$, where $\tau$ is the electron mean collision rate or the electron momentum scattering time), where, by definition, it accounts for temporally dispersive materials regardless of whether or not the material is lossy ($\Gamma_{d}$ accounts for loss, where $\Gamma_{d}=0$ describes lossless materials).
Then, conforming to the configuration modeled in Fig. 1, we use (40) and (41) for the permittivity expressions, where we let $\omega\rightarrow\omega_{0}$ for temporal region 1 and we Laplace transform the expressions ($\omega\rightarrow is$) for temporal region 2. We consider two cases. One is described in Table I, with a homogeneous dielectric space in temporal region 1 and a dielectric-plasma interface in temporal region 2. We also consider dielectric-plasma interfaces in both temporal regions, described in Table II.
TABLE I: Permittivity expressions for configuration of a homogeneous dielectric in temporal region 1 and a dielectric-plasma half-space in temporal region 2.
Temporal Region 1 Temporal Region 2
[…]
$\ \varepsilon_{r12}(s)=1+\frac{\omega_{p12}^{2}}{s^{2}+s\Gamma_{d12}}\\ \\ \end{array}$
TABLE II: Permittivity expressions for the configuration of a dielectric-plasma half-space in temporal region 1 and a different dielectric-plasma half-space in temporal region 2.
Temporal Region 1 Temporal Region 2
$\ \varepsilon_{r21}(\omega_{0})=\varepsilon_{r21}^{\prime}+i\frac{\sigma_{21}}{\omega_{0}\varepsilon_{0}}\\ \\ \end{array}$ $\ \varepsilon_{r22}(s)=\varepsilon_{r22}^{\prime}+\frac{\sigma_{22}}{s\varepsilon_{0}}\\ \\ \end{array}$
$\ \varepsilon_{r11}(\omega_{0})=1-\frac{\omega_{p11}^{2}}{\omega_{0}^{2}+i\omega_{0}\Gamma_{d11}}\\ \\ \end{array}$ $\ \varepsilon_{r12}(s)=1+\frac{\omega_{p12}^{2}}{s^{2}+s\Gamma_{d12}}\\ \\ \end{array}$
For the dielectric regions we use $\varepsilon_{r21}^{\prime}=\varepsilon_{r22}^{\prime}=1$ and $\sigma_{21}=\sigma_{22}=0.001\omega_{0}\varepsilon_{0}$. That is, since we need all materials to be dispersive, i.e., at least slightly lossy, we will consider a model of air in the limiting low-loss case. In the case of the plasma regions we will specify the parameters per each scenario investigated in the applicable results sections. The source excitation frequency and polarization that we use for all of the results are $f_{0}=\nicefrac{{\omega_{0}}}{{\left(2\pi\right)}}=15$ THz and $\mathbf{j}_{s}=\widehat{\mathbf{z}}j_{zs}$, respectively.
In Figs. 2 and 3, we first plot the sinusoidal steady state response in the frequency domain for an air-plasma half-space (no time boundary) to gain some insight into SPP behavior. For these plots we use the principal field (the direct source field in air; obtained using (27) in (38)) and the scattered field approximated as (strongly dominated by) the residue contribution [45], where the total field is the principal field plus the approximate scattered field.
Figure 2 shows a typical SPP response, where we see the SPP characteristic of confinement to the surface (exponential decay away from the surface in the $z$-direction) [46]. Additionally, we see the effects of decay with increasing dipole-interface separation, and loss as the SPPs propagate along the surface, in the $\rho$-direction. We also see these effects, for a fixed $\rho$, in Fig. 3, where we can see the transition from the propagation in air to SPP propagation as we approach the surface from above, in the $z$-direction. At large distances away from the interface, the principal and total fields are approximately the same (the principal field is the dominant contribution to the total field), but closer to the interface the SPP (contribution from residue) dominates the field.
Figure 2: Real part of the z-component of each component of the electric field, showing the principal field (the direct source field in air; obtained using (27) in (38)) and the scattered field approximated as (strongly dominated by) the residue contribution [45], in arbitrary units, as a function of the observation point height $z$ and the in-plane source-observation separation distance $\rho=\sqrt{\left(x-x^{\prime}\right)^{2}+\left(y-y^{\prime}\right)^{2}}$, where the source point height is $z^{\prime}=\nicefrac{{\lambda_{0}}}{{100}}$ and the permittivity configuration is an air-plasma half-space, where the plasma parameters are $\omega_{p12}=1.5\omega_{0}$, $\Gamma_{d12}=0.001\omega_{p12}$; $\varepsilon_{r12}(\omega_{0})=-1.25+i0.003$ (we are using the configuration in Table I, albeit, we are only using the configuration in temporal region 2 on its own (no time boundary, no temporal region 1) for the steady state case).
Figure 3: Real part of the z-component of each respective sinusoidal steady-state electric field, showing the principal field (the direct source field in air; obtained using (27) in (38)), the scattered field approximated as (strongly dominated by) the residue contribution [45], and the total field, where the total field is the principal field plus the approximate scattered field, as a function of the observation point height $z$ above the interface: (a) $z=2\lambda_{0}$, (b) $z=\lambda_{0}$, (c) $z=\nicefrac{{\lambda_{0}}}{{2}}$, (d) $z=\nicefrac{{\lambda_{0}}}{{3}}$, (e) $z=\nicefrac{{\lambda_{0}}}{{10}}$, (f) $z=\nicefrac{{\lambda_{0}}}{{100}}$, where the in-plane source-observation separation distance is $\rho=2.2\lambda_{0}$, the source point height is $z^{\prime}=\nicefrac{{\lambda_{0}}}{{100}}$, and the permittivity configuration is an air-plasma half-space, where the plasma is $\omega_{p12}=1.5\omega_{0}$, $\Gamma_{d12}=0.001\omega_{p12}$, $\varepsilon_{r12}=-1.25+i0.003$ (we are using the configuration in Table I, albeit, we are only using the configuration in temporal region 2 on its own (no time boundary, no temporal region 1) for the steady state case) for all cases.
III-A Dynamic SPP Formation
We now introduce a temporal interface and examine the field response versus time (normalized time defined as $\left(\nicefrac{{\nu_{phSPP2ss}}}{{\rho}}\right)t$, where $\nu_{phSPP2ss}=\nicefrac{{\omega_{0}}}{{\mathrm{Re}\left\{q_{SPP2ss}\left(\omega_{0}\right)\right\}}}$ is the phase velocity for the steady state SPP in temporal region 2), at various heights above the interface, to demonstrate dynamic SPP formation. We use the configuration in Table I, where the temporal region 2 plasma parameters are $\omega_{p12}=1.5\omega_{0}$, $\Gamma_{d12}=0.001\omega_{p12}$. In Fig. 4, before the time-boundary we plot the electric field in temporal region 1, $E_{z21}$, which is simply the usual sinusoidal response to a time-harmonic source. Starting at the time-boundary, when the media configuration changes from a homogeneous dielectric space to a dielectric-plasma interface (Table I) we have $E_{z22}$, forming a transient response. For reference, we also plot the sinusoidal steady-state response for the temporal region 2 media configuration, $E_{z22ss}$, which the transient field must tend to (for $t\gg\nicefrac{{\nu_{phSPP2ss}}}{{\rho}}$) if the response is stable. The fields are plotted for different observation point heights, compared to the wavelength $\lambda_{0}$, specified in the caption of Fig. 4. We obtain the fields, using the total Green’s function (the principal Green’s function obtained from (27) plus the scattered Green’s function obtained from (28)) in (38), where the scattered Green’s function is not approximated as its residue; it is obtained from the formal integration. We can see that at a height greater than a wavelength $\lambda_{0}$ the field in air (i.e., coupling to a continuum, as opposed to the SPP, which is a guided mode) dominates the total field response in temporal region 2, however, for heights less than a wavelength $\lambda_{0}$, the SPP starts to contribute to the total field. We can also see that the fields are continuous (the field in temporal region 1 (green) and the field in temporal region 2 (blue), which includes the transient response) at the temporal boundary $t=0$.
Figure 4: Real part of the electric field versus time for the steady state response in temporal region 1 ($E_{z21}$), the steady state response in temporal region 2 ($E_{z22ss}$), and the transient electric field in temporal region 2 ($E_{z22}$), as a function of the observation point height $z$ above the interface: (a) $z=2\lambda_{0}$, with additional plots, zoomed in around $t=0$ (temporal boundary), (b) $z=\nicefrac{{\lambda_{0}}}{{2}}$, with additional plots, zoomed in around $t=0$ (temporal boundary), (c) $z=\nicefrac{{\lambda_{0}}}{{3}}$, and (d) $z=\nicefrac{{\lambda_{0}}}{{5}}$, where the in-plane source-observation separation distance is $\rho=2.2\lambda_{0}$, the source point height is $z^{\prime}=\nicefrac{{\lambda_{0}}}{{100}}$, and we use the configuration in Table I, where the temporal region 2 plasma parameters are $\omega_{p12}=1.5\omega_{0}$, $\Gamma_{d12}=0.001\omega_{p12}$, $\varepsilon_{r12}=-1.25+i0.003$ for all cases. We can see that the fields are continuous (the field in temporal region 1 (green) and the field in temporal region 2 (blue), which includes the transient response) at the temporal boundary $t=0$. Note that $R=\left|\mathbf{r}-\mathbf{r}^{\prime}\right|=\sqrt{\rho^{2}+\left(z-z^{\prime}\right)^{2}}$, where $\rho=\sqrt{\left(x-x^{\prime}\right)^{2}+\left(y-y^{\prime}\right)^{2}}$.
As a whole, these plots demonstrate the dynamic SPP formation as the media is suddenly changed from air to an air-plasma half-space.
III-A1 Transient Period and Time to Steady State
We now consider the effect of different permittivity configurations (a change to the plasma material) in temporal region 2 on the SPP transient period and time to steady state. In this case, we use a source point height of $z^{\prime}=\nicefrac{{\lambda_{0}}}{{15}}$ and we use the configuration in Table I. In Fig. 5, where we have plots similar to the plots in Figs. 2 (the 3D plots) and 4 (the line plots), we can see that as we increase the plasma frequency the plasma permittivity becomes more negative (more like a metal), and the electrons in the plasma can more quickly respond to the excitation screening the field, resulting in a less confined SPP and longer SPP wavelengths. Furthermore, more reflection back into the air region can occur. Therefore, as the plasma permittivity becomes more negative, the transient response becomes faster, the SPP amplitude decreases (as we go from subfigure (a) to (d) in Fig. 5 we can see that the overall amplitude decreases as a result of more radiation into the air region), and steady state is approached much sooner.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 5: Plots similar to the plots in Figs. 2 (the 3D plots) and 4 (the line plots), where now we consider the effect of different permittivity configurations (the plasma material) in temporal region 2. We use the configuration in Table I, where we increase the plasma frequency $\omega_{p12}$: (a) $\omega_{p12}=1.5\omega_{0}$, $\varepsilon_{r12}=-1.25+i0.003$, (b) $\omega_{p12}=1.6\omega_{0}$, $\varepsilon_{r12}=-1.56+i0.004$, (c) $\omega_{p12}=1.7\omega_{0}$, $\varepsilon_{r12}=-1.89+i0.005$, and (d) $\omega_{p12}=2.5\omega_{0}$, $\varepsilon_{r12}=-5.25+i0.016$. In all cases, the source excitation frequency is $f_{0}=\nicefrac{{\omega_{0}}}{{\left(2\pi\right)}}=15$ THz, the damping frequency in the plasma material in temporal region 2 is $\Gamma_{d12}=0.001\omega_{p12}$, the in-plane source-observation separation distance is $\rho=2.2\lambda_{0}$, the source point height is $z^{\prime}=\nicefrac{{\lambda_{0}}}{{15}}$, and the observation point height is $z=\nicefrac{{\lambda_{0}}}{{15}}$. Note that $R=\left|\mathbf{r}-\mathbf{r}^{\prime}\right|=\sqrt{\rho^{2}+\left(z-z^{\prime}\right)^{2}}$, where $\rho=\sqrt{\left(x-x^{\prime}\right)^{2}+\left(y-y^{\prime}\right)^{2}}$.
Although not shown, as we increase the in-plane observation-source separation, the transient period is longer and it takes longer to approach steady state, since it takes more time for the SPP to travel farther; the system response time is longer.
Additionally, considering the effect of additional loss in the plasma material in temporal region 2 on the SPP response, at a fixed $\rho$, we observe that as we increase the loss in the plasma material the SPP quickly dampens out, which results in a shorter transient period and time to steady state.
III-B Interference of Slow and Fast Propagating SPPs
Here, we consider the interference of slow and fast propagating SPPs, using the configuration in Table II. The dipole source exists in both temporal regions. In temporal region 1 it excites an SPP that propagates through the temporal interface (that SPP, however, is no longer excited by the source once the interface has changed at the temporal boundary). In temporal region 2 the source excites a different SPP, according to the new material configuration. These two SPPs can interfere until the SPP originally excited in temporal region 1 dissipates. We use the SPP phase velocity to determine the SPP arrival time to the observation point, i.e., the time it takes to travel the in-plane source-observation separation distance, where $\nu_{phSPP1}=\nicefrac{{\omega_{0}}}{{\mathrm{Re}\left\{q_{SPP1}\left(\omega_{0}\right)\right\}}}$ and $\nu_{phSPP2ss}=\nicefrac{{\omega_{0}}}{{\mathrm{Re}\left\{q_{SPP2ss}\left(\omega_{0}\right)\right\}}}$. We can then determine the arrival time for SPPs created in temporal region 1 and the SPPs created in temporal region 2, starting from the material time change, i.e., $t=0$, as $t_{SPP1}=\nicefrac{{\rho}}{{\nu_{phSPP1}}}$ and $t_{SPP2}=\nicefrac{{\rho}}{{\nu_{phSPP2ss}}}$, respectively.
We consider two completely separate permittivity configuration scenarios, with an in-plane source-observation separation distance of $\rho=6\lambda_{0}$, where $\lambda_{0}=\nicefrac{{2\pi}}{{k_{0}}}\approx 20$ $\mu$m is vacuum wavelength. The first scenario is (a): a slow propagating SPP in temporal region 1, and a fast propagating SPP in temporal region 2. We then swap the configurations for temporal region 1 and 2, to see the opposite SPP response, where we then have (b): a fast propagating SPP in temporal region 1, and a slow propagating SPP in temporal region 2. Note that (a) and (b) are completely separate scenarios, where, in general, a faster propagating SPP corresponds to the plasma with a larger plasma frequency.
Then, for the corresponding SPP phase velocities of $\nu_{phSPP1}=1.8\times 10^{8}$ and $\nu_{phSPP2ss}=2.7\times 10^{8}$, the arrival times for scenario (a) are $t_{SPP1_{a}}=0.67$ ps and $t_{SPP2_{a}}=0.44$ ps, and for scenario (b) ($\nu_{phSPP1}=2.7\times 10^{8}$ and $\nu_{phSPP2ss}=1.8\times 10^{8}$), they are $t_{SPP1_{b}}=0.44$ ps and $t_{SPP2_{b}}=0.67$ ps, where the a and b subscripts correspond to the (a) and (b) scenarios that we are investigating, respectively. Those arrival times are shown in the plots in Fig. 6 (the green and red dashed lines). Note that when we change the media (the plasma) at $t=0$, the excitation of the SPPs for the plasma in temporal region 1 no longer occurs, instead the excitation of the SPPs for the plasma in temporal region 2 begins and continues. Since the source remains before and after the time change, any direct radiation from that will continue. In all cases it will take time for the fields to reach and pass the observation point. In the case of the SPP fields, the time it takes will depend on the SPP phase velocity.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 6: Real part of the electric field versus time plots regarding interference of slow and fast propagating SPPs. We use the configuration in Table II, where the permittivity configuration for the plasma in temporal region 1 and 2 is: (a) slow propagating SPP: $\omega_{p11}=1.6\omega_{0}$, $\Gamma_{d11}=0.001\omega_{p11}$, $\varepsilon_{r11}=-1.56+i0.004$ and fast propagating SPP: $\omega_{p12}=2.5\omega_{0}$, $\Gamma_{d12}=0.001\omega_{p12}$, $\varepsilon_{r12}=-5.25+i0.016$, respectively, and (b) fast propagating SPP: $\omega_{p11}=2.5\omega_{0}$, $\Gamma_{d11}=0.001\omega_{p11}$, $\varepsilon_{r11}=-5.25+i0.016$ and slow propagating SPP: $\omega_{p12}=1.6\omega_{0}$, $\Gamma_{d12}=0.001\omega_{p12}$, $\varepsilon_{r12}=-1.56+i0.004$, respectively, where the additional plots are zoomed in around the time range where the slow and fast propagating SPPs are potentially able to interfere given their arrival time (green dashed line is for the SPP from temporal region 1, red dashed line is for the SPP formed in temporal region 2) to the observation point. In all cases, the in-plane source-observation separation distance is $\rho=6\lambda_{0}$, the source point height is $z^{\prime}=\nicefrac{{\lambda_{0}}}{{15}}$, and the observation point height is $z=\nicefrac{{\lambda_{0}}}{{15}}$. Note that $R=\left|\mathbf{r}-\mathbf{r}^{\prime}\right|=\sqrt{\rho^{2}+\left(z-z^{\prime}\right)^{2}}$, where $\rho=\sqrt{\left(x-x^{\prime}\right)^{2}+\left(y-y^{\prime}\right)^{2}}$.
Therefore, the SPP from temporal region 1, which is present at the observation point before the time change, will continue to be present at (i.e., propagating past) the observation point until the last part of the wave, which ceases to be excited after $t=0$, reaches (passes by) the observation point, after which it will cease to exist (eventually dying out). The SPP created in temporal region 2, starting at the time change at $t=0$, will also take time to reach the observation point, where it will have its own arrival time depending on its phase velocity, after which it will continue to be excited by the source and propagate. With all of this taken into consideration we can investigate the interference of slow and fast propagating SPPs based on their arrival times.
In the case of a slow propagating SPP from temporal region 1 and a fast propagating SPP in temporal region 2 (Fig. 6a), the slow propagating SPP (from temporal region 1) has been continuously passing by the observation point (has always been there). It takes a relatively long time to stop passing the observation point (0.67 ps). Meanwhile, starting at 0.44 ps, the fast propagating SPP, created (starting) at $t=0$, reaches the observation point, and since the slow propagating SPP is still there, the two interfere. After 0.67 ps the slow propagating SPP is gone, and interference ends, which is what we see at 0.67 ps (the green dashed line) in the plot in Fig. 6a. Given this we see that we can interfere a slow and fast propagating SPP in a known time frame to induce constructive interference between the SPPs and achieve a larger amplitude.
We also see interference occurring before this time, after the fast propagating SPP (from temporal region 2) arrival time and before it. This interference is occurring between the forward and backward waves (from the SPPs and direct radiation), which have shifted frequencies relative to the those waves in temporal region 1, that result due to the momentum conservation at the temporal boundary. Up to this point we haven’t discussed this much, however, we can see it occurring in some of the other previous plots, where it is more pronounced, e.g., in Fig. 4. Based on these results we can see that it is possible to enhance the SPP field intensity by interfering slow and fast propagating SPPs.
In Fig. 6b, the case of a fast propagating SPP from temporal region 1 and a slow propagating SPP in temporal region 2, the fast propagating SPP (from temporal region 1) stops passing the observation point after 0.44 ps, before the slow propagating SPP, created (starting) at $t=0$, reaches the observation point (after 0.67 ps). Therefore, the SPPs are unable to interfere with each other, which is what we see (there is no interference) around that time frame in the plot in Fig. 6b.
IV Conclusion
We established the theoretical framework for the specific configuration of a time-varying media system that supports SPPs excited by a dipole excitation (electric dipole point source), where the media configuration suddenly changes at a temporal boundary. Such a system allows for the analysis of the interaction among dipole excitations, where we have the ability to consider the interactions at the moment of SPP creation. Using this framework, we then demonstrated dynamic SPP formation and interference of slow and fast propagating SPPs, for the case of isotropic media and one temporal boundary. This provided insight into how SPPs respond in a time-varying media system and, in the case of the interference of slow and fast propagating SPPs, demonstrated that we can induce some constructive interference between SPPs in a known time frame. There are numerous avenues that warrant further research in the area of time-varying media, where further work is needed to incorporate the case of anisotropic media and multi-temporal boundaries into the framework already established here for this type of system.